9.yto the power 4 +20 = 29 y square
Answer by mananth(16946)

9y^4+20=29y^2
9y^4-29Y^2+20=0
9y^4-20y^2-9y^2+20=0
y^2(9y^2-20)-1(9y^2-20)
(9y^2-20)(y^2-1)
(9y^2-20)(y+1)(y-1)
(3y+sqrt20)(3y-sqrt20)(y+1)(y-1)
...
m.ananth@hotmail.ca